Gelbart Company manufactures gas grills. Fixed costs amount to $20,152,800 per year. Variable costs per gas grill are $360, and the average price per gas grill is $900.

Required:

1. How many gas grills must Gelbart Company sell to break even?
gas grills

2. If Gelbart Company sells 40,310 gas grills in a year, what is the operating income?
$

3. If Gelbart Company’s variable costs increase to $378 per grill while the price and fixed costs remain unchanged, what is the new break-even point? If required, round your answer to the nearest whole number.
gas grills
Answer #1

1. Break even point = Fixed costs / Contribution margin per unit

= $20,152,800 / ($900 - 360) = 37320 gas grills

2. Operating income = Sales - Variable cost - Fixed cost

= 900*40310 - 360*40310 - 20152800

= $1,614,600

3. Break even point = Fixed costs / Contribution margin per unit

= $20,152,800 / ($900 - 378) = 38,607 gas grills